# CaRe-BN: Precise Moving Statistics for Stabilizing Spiking Neural Networks in Reinforcement Learning

### v1

@@ Abstract
Spiking Neural Networks (SNNs) offer low-latency and energy-efficient decision-making on neuromorphic hardware by mimicking the event-driven dynamics of biological neurons. However, due to the discrete and non-differentiable nature of spikes, directly trained SNNs rely heavily on Batch Normalization (BN) to stabilize gradient updates. In online Reinforcement Learning (RL), imprecise BN statistics hinder exploitation, resulting in slower convergence and suboptimal policies. This challenge limits the adoption of SNNs for energy-efficient control on resource-constrained devices. To overcome this, we propose Confidence-adaptive and Re-calibration Batch Normalization (CaRe-BN), which introduces (i) a confidence-guided adaptive update strategy for BN statistics and (ii) a re-calibration mechanism to align distributions. By providing more accurate normalization, CaRe-BN stabilizes SNN optimization without disrupting the RL training process. Importantly, CaRe-BN does not alter inference, thus preserving the energy efficiency of SNNs in deployment. Extensive experiments on continuous control benchmarks demonstrate that CaRe-BN improves SNN performance by up to $22.6\%$ across different spiking neuron models and RL algorithms. Remarkably, SNNs equipped with CaRe-BN even surpass their ANN counterparts by $5.9\%$. These results highlight a new direction for BN techniques tailored to RL, paving the way for neuromorphic agents that are both efficient and high-performing.
@@ 1 Introduction
In parallel, Reinforcement Learning (RL) has achieved remarkable success across a wide range of domains (Mnih et al., 2015b; Lillicrap et al., 2015; Haarnoja et al., 2018). Among these, continuous control tasks have received significant attention due to their alignment with real-world scenarios and their strong connection to embodied AI and robotic applications (Kober et al., 2013; Gu et al., 2017; Brunke et al., 2022). Combining the strengths of SNNs with RL (SNN-RL) offers the potential to train agents that not only learn complex behaviors but also execute them with extremely low energy consumption (Yamazaki et al., 2022). This makes SNN-RL particularly appealing for robotics and autonomous systems deployed on resource-constrained edge devices.
However, training SNNs is challenging. Due to their discrete dynamics and non-differentiable activation functions, SNNs are heavily dependent on Batch Normalization (BN) (Ioffe & Szegedy, 2015) to mitigate gradient vanishing or explosion. BN stabilizes optimization by normalizing activations with moving estimates of the mean and variance, thereby alleviating temporal covariate shifts in SNNs and enabling state-of-the-art performance (Duan et al., 2022; Jiang et al., 2024).
@@
Figure 1: Real and estimated input activation distributions in BN layers. Distributions change rapidly in (a) and (c), while remaining stable in (b) and (d).
@@
We evaluate CaRe-BN on various continuous control tasks from MuJoCo (Todorov et al., 2012; Todorov, 2014b). The results show that CaRe-BN not only resolves the issue of imprecise BN statistics but also accelerates training and achieves state-of-the-art performance. Remarkably, SNN-based agents equipped with CaRe-BN even outperform their ANN counterparts by $\mathbf{5.9\%}$, without requiring complex neuron dynamics or specialized RL frameworks.
@@ 2.1 Batch Normalization in Spiking Neural Networks
Batch Normalization (BN) was originally proposed for ANNs to mitigate internal covariate shift during training (Ioffe & Szegedy, 2015), thereby accelerating convergence and improving performance (Santurkar et al., 2018). To address unstable training in SNNs, several extensions of BN have been developed (Zheng et al., 2021; Duan et al., 2022; Kim & Panda, 2021; Jiang et al., 2024). While these methods are effective in supervised tasks, they are designed under the assumption of static training–inference distributions. This assumption is violated in online RL, where distributions shift continually as the agent interacts with the environment, making these BN variants ill-suited for SNN-RL.
@@ 2.2 Spiking Neural Networks in Reinforcement Learning
Early work in SNN-RL primarily relied on synaptic plasticity rules, particularly reward-modulated Spike-Timing-Dependent Plasticity (R-STDP) and its variants (Florian, 2007; Frémaux & Gerstner, 2016; Gerstner et al., 2018; Frémaux et al., 2013; Yang et al., 2024). Another research direction focused on ANN-to-SNN conversion: Patel et al. (2019); Tan et al. (2021); Kumar et al. (2025) converted Deep Q-Networks (DQNs) (Mnih, 2013; Mnih et al., 2015a) into SNNs. To enable direct gradient-based training, Liu et al. (2022); Chen et al. (2022); Qin et al. (2022); Sun et al. (2022) applied Spatio-Temporal Backpropagation (STBP) (Wu et al., 2018) to train DQNs, while Bellec et al. (2020) introduced e-prop with eligibility traces to train policy networks using policy gradient methods (Sutton et al., 1999).
For continuous control tasks, hybrid frameworks have been extensively explored (Tang et al., 2020; 2021; Zhang et al., 2022; Chen et al., 2024a; Zhang et al., 2024; Ding et al., 2022; Chen et al., 2024b; Xu et al., 2025). These approaches typically employ a Spiking Actor Network (SAN) co-trained with a deep ANN critic in the Actor–Critic framework (Konda & Tsitsiklis, 1999). However, none of these methods address the challenge of normalization in SNN-based RL. The absence of proper normalization often leads to unstable updates, slower convergence, or even divergence during training.
@@ 3.2 Reinforcement Learning
Reinforcement Learning (RL) is a framework in which an agent learns to maximize cumulative rewards by interacting with an environment. The agent maps states (or observations) to actions, with the learning loop consisting of two steps: (i) the agent selects an action, receives a reward, and transitions to the next state; and (ii) the agent updates its policy by sampling mini-batches of past experiences.
@@ 3.3 Batch Normalization
Batch Normalization (BN) (Ioffe & Szegedy, 2015) is a widely used technique to stabilize and accelerate the training of deep neural networks. Given an activation $x_{i}\in\mathbb{R}^{d}$ at iteration $i$, BN normalizes it using the mean and variance computed over a mini-batch $\mathcal{B}=\{x_{i}^{1},\dots,x_{i}^{N}\}$:
@@ 4.1 Issues in Approximating Moving Statistics
It is worth noting that most conventional ANN-based RL algorithms do not employ BN (Lillicrap, 2015; Sutton & Barto, 2018), as shallow ANNs can often learn stable representations without normalization. In contrast, BN is indispensable for stabilizing SNNs training. Therefore, addressing this issue is particularly critical for SNN-based RL.
@@ 4.2 Confidence-adaptive Update of BN Statistics (Ca-BN)
Let $(\mu_{i},\sigma_{i}^{2})$ and $(\hat{\mu}_{i\mid i-1},\hat{\sigma}^{2}_{i\mid i-1})$ be two unbiased estimators of the population parameters $(\mu_{i}^{*},{\sigma_{i}^{*}}^{2})$. The optimal linear estimator that minimizes MSE is
@@
where $K^{\mu}_{i}$ and $K^{\sigma}_{i}$ are confidence-guided adaptive weights, and $\mathbb{D}(\cdot)$ denotes estimation variance (the reciprocal of confidence).
@@
Since $\mu_{i}$ and $\sigma_{i}^{2}$ are unbiased estimators of $\mu^{*}_{i}$ and ${\sigma_{i}^{*}}^{2}$, we approximate
@@
Confidence of previous estimates. Since the true distribution drift is unknown, we approximate
$\hat{\mu}_{i\mid i-1}=\hat{\mu}_{i-1},\qquad\hat{\sigma}^{2}_{i\mid i-1}=\hat{\sigma}^{2}_{i-1},$ (10)
@@
where $D_{i}^{\mu},D_{i}^{\sigma}$ are recursively smoothed statistics with momentum $\alpha$:
$D^{\mu}_{i}\leftarrow(1-\alpha)D^{\mu}_{i-1}+\alpha(\mu_{i}-\hat{\mu}_{i-1})^{2},\qquad D^{\sigma}_{i}\leftarrow(1-\alpha)D^{\sigma}_{i-1}+\alpha(\sigma_{i}^{2}-\hat{\sigma}^{2}_{i-1})^{2}.$ (12)
Combining Eqs. 512, we obtain the confidence-adaptive update scheme. When distributional shifts are rapid, $D_{i}^{\mu}$ and $D_{i}^{\sigma}$ grow large, increasing $K^{\mu}_{i}$ and $K^{\sigma}_{i}$ and accelerating adaptation. Conversely, when statistics are stable, these terms shrink, lowering $K^{\mu}_{i}$ and $K^{\sigma}_{i}$ and reducing noise from small mini-batches.
@@ 4.3 Re-calibration Mechanism of BN Statistics (Re-BN)
While the confidence-adaptive update provides online estimates of BN statistics during training, these estimates may still drift from the true population values due to stochastic mini-batch noise. The most accurate approach would be to recompute exact statistics by forward-propagating the entire dataset after each update (Wu & Johnson, 2021). However, this is computationally infeasible in RL, as it would require processing millions of samples at every step.
@@ 4.4 Integrating with RL
It is important to note that the inference procedure of CaRe-BN remains identical to that of conventional BN. Consequently, the CaRe-BN layer is seamlessly fused into synaptic weights, introducing no additional inference overhead during deployment.
@@ 5 Experiments
5.1 Experiment Setup
We evaluate our CaRe-BN on five standard continuous control benchmarks from the MuJoCo suite (Todorov et al., 2012; Todorov, 2014b) within the OpenAI Gymnasium (Brockman, 2016; Towers et al., 2024), including InvertedDoublePendulum (IDP) (Todorov, 2014a), Ant (Schulman et al., 2015), HalfCheetah (Wawrzyński, 2009), Hopper (Erez et al., 2012), and Walker2d. All environments use default settings, and performance is evaluated by averaging the rewards over $10$ trials.
To assess the effectiveness and adaptability of CaRe-BN, we conduct experiments with different spiking neuron models and RL algorithms. For spiking neurons, we test the Leaky Integrate-and-Fire (LIF) (Gerstner & Kistler, 2002) and Current-based LIF (CLIF) (Tang et al., 2021), with the dynamics detailed in Appendix. For RL algorithms, we employ Deep Deterministic Policy Gradient (DDPG) (Lillicrap, 2015) and Twin Delayed DDPG (TD3) (Fujimoto et al., 2018). All SNNs are simulated with a time step of $5$, and additional training details are provided in Appendix.
For a fair comparison, we maintain consistent hyperparameters across all models and algorithms, with full settings provided in Appendix. Each algorithm is trained using $5$ random seeds, and we report the averaged results.
@@ 5.2 More Precise BN Statistics Lead to Better Exploration
A key aspect of online RL is that the quality of exploration significantly influences subsequent policy updates. Figure 3 compares the exploration performance of standard BN with our proposed CaRe-BN. Throughout the training process, CaRe-BN consistently achieves higher exploration returns.
Figure 3: Exploration performance of BN and CaRe-BN with CLIF neurons and the TD3 algorithm across five MuJoCo tasks. Shaded areas represent half a standard deviation across five random seeds. Curves are uniformly smoothed for visual clarity.
Since CaRe-BN does not directly modify the gradient update process, the observed improvement in exploration performance is solely due to its more precise estimation of BN statistics. This leads to better exploration policies, which in turn generate higher-quality trajectories for updating the agent. As a result, CaRe-BN forms a positive feedback loop: improved statistics $\rightarrow$ better exploration $\rightarrow$ higher-quality experiences $\rightarrow$ better policy.

### v2

@@ Abstract
Spiking Neural Networks (SNNs) offer low-latency and energy-efficient decision-making on neuromorphic hardware by mimicking the event-driven dynamics of biological neurons. However, the discrete and non-differentiable nature of spikes leads to unstable gradient propagation in directly trained SNNs, making Batch Normalization (BN) an important component for stabilizing training. In online Reinforcement Learning (RL), imprecise BN statistics hinder exploitation, resulting in slower convergence and suboptimal policies. While Artificial Neural Networks (ANNs) can often omit BN, SNNs critically depend on it, limiting the adoption of SNNs for energy-efficient control on resource-constrained devices. To overcome this, we propose Confidence-adaptive and Re-calibration Batch Normalization (CaRe-BN), which introduces (i) a confidence-guided adaptive update strategy for BN statistics and (ii) a re-calibration mechanism to align distributions. By providing more accurate normalization, CaRe-BN stabilizes SNN optimization without disrupting the RL training process. Importantly, CaRe-BN does not alter inference, thus preserving the energy efficiency of SNNs in deployment. Extensive experiments on both discrete and continuous control benchmarks demonstrate that CaRe-BN improves SNN performance by up to $22.6\%$ across different spiking neuron models and RL algorithms. Remarkably, SNNs equipped with CaRe-BN even surpass their ANN counterparts by $5.9\%$. These results highlight a new direction for BN techniques tailored to RL, paving the way for neuromorphic agents that are both efficient and high-performing. Code is available at https://github.com/xuzijie32/CaRe-BN.
@@ 1 Introduction
In parallel, Reinforcement Learning (RL) has achieved remarkable success across a wide range of domains (Mnih et al., 2015; Lillicrap et al., 2015; Haarnoja et al., 2018a). Among these, complex control tasks have received significant attention due to their alignment with real-world scenarios and their strong connection to embodied AI and robotic applications (Kober et al., 2013; Gu et al., 2017; Brunke et al., 2022). Combining the strengths of SNNs with RL (SNN-RL) offers the potential to train agents that not only learn complex behaviors but also execute them with extremely low energy consumption (Yamazaki et al., 2022). This makes SNN-RL particularly appealing for robotics and autonomous systems deployed on resource-constrained edge devices.
However, training SNNs is challenging. Due to the discrete spike dynamics and the reliance on surrogate gradients to approximate the backward pass, directly trained SNNs often suffer from unstable gradient propagation, including vanishing or exploding gradients (Zheng et al., 2021). Batch Normalization (BN) (Ioffe and Szegedy, 2015) plays a crucial role in stabilizing SNN training by regulating activation statistics and improving gradient flow, mitigates such instability and contributes to state-of-the-art performance (Duan et al., 2022; Jiang et al., 2024).
@@
This issue is especially critical for SNNs. Traditional online RL algorithms usually remove BN layers in their networks (Sutton and Barto, 2018; Fujimoto et al., 2018; Haarnoja et al., 2017; Schulman et al., 2017). Unlike ANNs that can train stably without BN, SNNs rely heavily on normalization to stabilize membrane potentials and surrogate-gradient backpropagation. Removing BN from SNN-based RL leads to severe gradient instability and substantial performance degradation.
Figure 1: Real and estimated input activation distributions in BN layers. Between each gradient update iterations, distributions change rapidly in (a) and (c), while remaining stable in (b) and (d).
@@
We evaluate CaRe-BN on a variety of control tasks, including the Atari benchmark (Bellemare et al., 2013; Machado et al., 2018) for discrete action spaces and the MuJoCo suite (Todorov et al., 2012; Todorov, 2014b) for continuous control. The results show that CaRe-BN not only resolves the issue of imprecise BN statistics but also accelerates training and achieves state-of-the-art performance. Remarkably, SNN-based agents equipped with CaRe-BN even outperform their ANN counterparts by $\mathbf{5.9\%}$, without requiring complex neuron dynamics or specialized RL frameworks.
@@ 2.1 Batch Normalization in Spiking Neural Networks
Batch Normalization (BN) was originally proposed for ANNs to mitigate internal covariate shift during training (Ioffe and Szegedy, 2015), thereby accelerating convergence and improving performance (Santurkar et al., 2018). To address unstable training in SNNs, several extensions of BN have been developed (Zheng et al., 2021; Duan et al., 2022; Kim and Panda, 2021; Jiang et al., 2024). While these methods are effective in supervised tasks, they are designed under the assumption of static distributions. This assumption is violated in online RL, where distributions shift continually as the agent interacts with the environment, making these BN variants ill-suited for SNN-RL.
@@ 2.2 Spiking Neural Networks in Reinforcement Learning
Early work in SNN-RL primarily relied on synaptic plasticity rules, particularly reward-modulated Spike-Timing-Dependent Plasticity (R-STDP) and its variants (Florian, 2007; Frémaux and Gerstner, 2016; Gerstner et al., 2018; Frémaux et al., 2013; Yang et al., 2024). Another research direction focused on ANN-to-SNN conversion: while Patel et al. (2019); Tan et al. (2021); Kumar et al. (2025) converted Deep Q-Networks (DQNs) (Mnih, 2013; Mnih et al., 2015) into SNNs for discrete control, recent work by Xu et al. (2026) has greatly reduced conversion errors in continuous control. To enable direct gradient-based training, Liu et al. (2022); Chen et al. (2022); Qin et al. (2022); Sun et al. (2022) applied Spatio-Temporal Backpropagation (STBP) (Wu et al., 2018) to train DQNs, while Bellec et al. (2020) introduced e-prop with eligibility traces to train policy networks using policy gradient methods (Sutton et al., 1999).
For continuous control tasks, hybrid frameworks have been extensively explored (Tang et al., 2020; 2021; Zhang et al., 2022; Chen et al., 2024a; Zhang et al., 2024; Ding et al., 2022; Chen et al., 2024b). These approaches typically employ a Spiking Actor Network (SAN) co-trained with a deep ANN critic in the Actor–Critic framework (Konda and Tsitsiklis, 1999). Empowered by the recent proxy target framework (Xu et al., 2025), simple SNNs are now capable of outperforming their ANN counterparts. However, none of these methods address the challenge of normalization in SNN-based RL. The absence of proper normalization often leads to unstable updates and slower convergence.
@@ 3.2 Reinforcement Learning
RL is a framework in which an agent learns to maximize cumulative rewards by interacting with an environment. The agent maps states (or observations) to actions, with the learning loop consisting of two steps: (i) the agent selects an action, receives a reward, and transitions to the next state; and (ii) the agent updates its policy by sampling mini-batches of past experiences.
@@ 3.3 Batch Normalization
Batch Normalization (BN) (Ioffe and Szegedy, 2015) is a widely used technique to stabilize and accelerate the training of deep neural networks. Given an activation $x_{i}\in\mathbb{R}^{d}$ at iteration $i$, BN normalizes it using the mean and variance computed over a mini-batch $\mathcal{B}=\{x_{i}^{1},\dots,x_{i}^{N}\}$:
@@ 4.1 Issues in Approximating Moving Statistics
It is worth noting that most conventional ANN-based RL algorithms do not employ BN (Lillicrap, 2015; Sutton and Barto, 2018), as shallow ANNs can often learn stable representations without normalization. In contrast, BN is indispensable for stabilizing SNNs training. Therefore, addressing this issue is particularly critical for SNN-based RL.
@@ 4.2 Confidence-adaptive Update of BN Statistics (Ca-BN)
Let $(\mu_{i},\sigma_{i}^{2})$ and $(\hat{\mu}_{i\mid i-1},\hat{\sigma}^{2}_{i\mid i-1})$ be two unbiased estimators of the population parameters $(\mu_{i}^{*},{\sigma_{i}^{*}}^{2})$. Taking them as random variables, the optimal linear estimator is
@@
where $K^{\mu}_{i}$ and $K^{\sigma}_{i}$ are confidence-guided adaptive weights, and $\mathbb{D}(\cdot)$ denotes generalized variance [^1] The confidence is defined as the inverse of the generalized variance: $\text{confidence score}=\frac{1}{\mathbb{D}}$..
@@
Since $\mu^{*}_{i}$ and ${\sigma_{i}^{*}}^{2}$ are unknown, we adopt the common approximation using $\mu_{i}$ and $\sigma_{i}^{2}$, thus:
@@
Confidence of previous estimates. Since the true statistics $\mu_{i}^{*}$ and ${\sigma_{i}^{*}}^{2}$ are unknown, direct computation of $\mathbb{D}(\mu_{i}^{*}-\hat{\mu}_{i\mid i-1})$ and $\mathbb{D}({\sigma_{i}^{*}}^{2}-\hat{\sigma}^{2}_{i\mid i-1})$ is infeasible. To approximate them, we view the minibatch statistics $\mu_{i}$ and $\sigma_{i}^{2}$ as a stochastic sample drawn from the unknown hypothetical distributions induced by $\mu_{i}^{*}$ and ${\sigma_{i}^{*}}^{2}$. Thus, the squared deviations $(\mu_{i}-\hat{\mu}_{i\mid i-1}))^{2}$ and $(\sigma_{i}^{2}-\hat{\sigma}^{2}_{i\mid i-1}))^{2}$ serve as unbiased but noisy probes of $\mathbb{D}(\mu_{i}^{*}-\hat{\mu}_{i\mid i-1})$ and $\mathbb{D}({\sigma_{i}^{*}}^{2}-\hat{\sigma}^{2}_{i\mid i-1})$.
Because these single-minibatch estimates exhibit high variance, we maintain smoothed recursive estimators updated using an exponential moving average with momentum parameter $\alpha$:
@@
$D^{\mu}_{i}\leftarrow(1-\alpha)D^{\mu}_{i-1}+\alpha(\mu_{i}-\hat{\mu}_{i\mid i-1}))^{2},\qquad D^{\sigma}_{i}\leftarrow(1-\alpha)D^{\sigma}_{i-1}+\alpha(\sigma_{i}^{2}-\hat{\sigma}^{2}_{i\mid i-1}))^{2}.$ (11)
Combining Eqs. 511, we obtain the confidence-adaptive update scheme [^2] As BN statistics fluctuate without monotonic trends, we define $\hat{\mu}_{i\mid i-1}=\mu_{i-1}$ and $\hat{\sigma}^{2}_{i\mid i-1}=\sigma^{2}_{i-1}$.. When distributional shifts are rapid, $D_{i}^{\mu}$ and $D_{i}^{\sigma}$ grow large, increasing $K^{\mu}_{i}$ and $K^{\sigma}_{i}$ and accelerating adaptation. Conversely, when statistics are stable, these terms shrink, lowering $K^{\mu}_{i}$ and $K^{\sigma}_{i}$ and reducing noise from small mini-batches.
@@ 4.3 Re-calibration Mechanism of BN Statistics (Re-BN)
While the confidence-adaptive update provides online estimates of BN statistics during training, these estimates may still drift from the true population values due to stochastic mini-batch noise. The most accurate approach would be to recompute exact statistics by forward-propagating the entire dataset after each update (Wu and Johnson, 2021). However, this is computationally infeasible in RL, as it would require processing millions of samples at every step.
@@ 4.4 Integrating with RL
It is important to note that the inference procedure of CaRe-BN remains identical to that of conventional BN. At inference time, the CaRe-BN layer is seamlessly fused into synaptic weights, introducing no additional inference overhead during deployment.
@@ 5 Experiments
5.1 Experimental Setup
We evaluate CaRe-BN on RL tasks covering both discrete and continuous action spaces. All environments use default settings, and performance is evaluated by averaging the rewards in $10$ trials.
For discrete action spaces, we consider four widely used Atari 2600 games from the Arcade Learning Environment (ALE) (Bellemare et al., 2013; Machado et al., 2018): Pong, Breakout, SpaceInvaders, Freeway, and Seaquest. We adopt a deep Q-learning framework (Mnih et al., 2015) and train a deep Spiking Q-Network (Liu et al., 2022) that receives RAM-based observations and outputs state-action values.
For continuous control, we evaluate on five standard MuJoCo benchmarks (Todorov et al., 2012; Todorov, 2014b) provided in the OpenAI Gymnasium suite (Brockman, 2016; Towers et al., 2024): InvertedDoublePendulum (IDP) (Todorov, 2014a), Ant (Schulman et al., 2015), HalfCheetah (Wawrzyński, 2009), Hopper (Erez et al., 2012), and Walker2d. We employ a hybrid framework in which a spiking actor network is co-trained with a deep critic network using several RL algorithms, including Deep Deterministic Policy Gradient (DDPG) (Lillicrap, 2015), Twin Delayed DDPG (TD3) (Fujimoto et al., 2018), and Soft Actor-Critic (SAC) (Haarnoja et al., 2018b).
To evaluate the generality of CaRe-BN, we experiment with multiple spiking neuron models: the Leaky Integrate-and-Fire (LIF) neuron (Gerstner and Kistler, 2002), the Current-based LIF (CLIF) neuron (Tang et al., 2021), and the Dynamic Neuron (DN) model (Zhang et al., 2022), with detailed dynamics provided in the Appendix. All SNN agents are trained via Spatio-Temporal Backpropagation (STBP) (Wu et al., 2018), with the CaRe-BN module inserted between every pair of adjacent layers. For fair comparison, all models share the same hyperparameters, fully listed in the Appendix.
During each RL environment step, the SNN agent performs a single forward inference composed of $5$ simulation time steps, after which all neuron states are reset.
@@ 5.2 More Precise BN Statistics Lead to Better Exploration
Figure 3: Wasserstein distance between estimated BN statistics and the true distribution across layers, measured with CLIF neurons and the TD3 algorithm in the InvertedDoublePendulum-v4 environment. Shaded areas denote half a standard deviation over five runs. Curves are uniformly smoothed for visual clarity.
In online RL, the quality of exploration directly affects subsequent policy updates. As discussed in Section 4.1, traditional BN methods struggle to maintain accurate moving statistics, which can lead to suboptimal exploration behavior.
To quantify this effect, we compute the Wasserstein distance between the true feature distribution and the Gaussian distribution estimated by BN. Figure 3 shows that CaRe-BN consistently reduces this discrepancy across all layers throughout training, producing more precise normalization.
Figure 4: Exploration returns of BN and CaRe-BN with CLIF neurons and the TD3 algorithm across five MuJoCo tasks. Shaded areas represent half a standard deviation across five random seeds. Curves are uniformly smoothed for visual clarity.
The impact of improved statistics is reflected in exploration performance. As shown in Figure 4, CaRe-BN consistently achieves higher exploration returns. Since CaRe-BN does not directly modify the gradient update process, the observed improvement in exploration performance is solely due to its more precise estimation of BN statistics. This leads to better exploration policies, which in turn generate higher-quality trajectories for updating the agent. As a result, CaRe-BN forms a positive feedback loop: improved statistics $\rightarrow$ better exploration $\rightarrow$ higher-quality experiences $\rightarrow$ better policy.